OpenStudy (anonymous):

Francis delivers 440 pamphlets in 2 hours and earns $5 per 100 pamphlets delivered. How long will it take Francis to earn $55?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

For earning $55 he will have to deliver : \(11 \times 1100 = 1100\) In 2 hours he delivers : 440 How much time will it take to deliver 1100 pamphlets, that we have to find..

OpenStudy (anonymous):

440 pamphlets can be delivered in : 2 hours: So: 1100 pamphlets will be delivered in : \(\frac{2}{440} \times 1100 = 5\). So in my view 5 hours it will take..

OpenStudy (anonymous):

ur first part shouldn't it be 11 x 100 = 1100 not 11 x 1100 = 1100?

OpenStudy (anonymous):

Oh my mistake.. Wait a minute..

OpenStudy (anonymous):

That will be 11 times 100 = 1100..

OpenStudy (anonymous):

and wat does the 11 represent?

OpenStudy (anonymous):

For earning $55 he will have to deliver : \(11 \times 100 = 1100\)..

OpenStudy (anonymous):

Let me explain..

OpenStudy (anonymous):

5 dollars he earns for : 100 pamphlets 1 dollar he earns for: \(\frac{100}{5} = 20\) pamphlets So 55 dollars he will earn for : \(20 \times 55 = 1100\)

OpenStudy (anonymous):

Getting now ??

OpenStudy (anonymous):

So then you have to find how much will it take to deliver 1100 pamphlets.. Are you getting or not @JayDS

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

and why is it 2/440 for the second part?

OpenStudy (anonymous):

See: 440 pamphlets can be delivered in : 2 hours So if you want to find delivery of 1 pamphlet you will divide both sides by 440.. So: \[\frac{440}{440} \; pamphlets \; In \; \frac{2}{440} \quad hours\] Getting ??

OpenStudy (anonymous):

So: 1 pamphlet can be delivered in : \(\frac{2}{440}\) and for 1100 you will multiply both the sides by 1100.. \(1 \times 1100\) pamphlets can be delivered in : \(\frac{2}{440} \times 1100 = 5\) Now is it clear to you ?

OpenStudy (anonymous):

the 2/440 is still a bit confusing.

OpenStudy (anonymous):

440 pamphlets can be delivered in : 2 hours this much is clear to you or not ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Write it like simple equation: 440x = 2y ( x represents pamphlets and y represents hours) Now to find x what will you do ??

OpenStudy (anonymous):

isolate x so x = 2y/440

OpenStudy (anonymous):

That is what I did there.. So I can write it as: 1x = 2/440 y Now clear to you ??

OpenStudy (anonymous):

replace here x any y: 1 pamphlet = 2/440 hours..

OpenStudy (anonymous):

ok, I think so, I will have to write all of it in sequence first.

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

you can help others if you want to, and thanks

OpenStudy (anonymous):

Okay.. Take your time..